Reports 1-1 of 1 Clear search Modify search
AdV-DET (Optical benches layout)
tacca, berni - 17:15 Saturday 05 June 2021 (52050) Print this report
Comment to SNEB upgrade for 40 W of input power (51979)

The implementation of the new SNEB configuration has been completed:

  • the North cavity has been locked on B7 in air;
  • the beam has been almost centered on the camera and the photodiodes moving both the bench and the picomotors;
  • the North cavity has been locked on B4;
  • the beam has been roughly centered on the quadrant B7_QD2: the beam in not clippend on M3, light is arriving on the sensors, it was difficult to understand the exact meaning of the B7_QD2 signals (to be further checked);
  • the beam reflected by M3 has been centered on the new beam dump installed last Monday;
  • the beam reflected by the quadrant has been sent on the corresponding beam dump;
  • the balancing of the bench has been checked;
  • the microtower has been closed and pumped up to a decent level of vacuum (slighlty lower than 1 mbar) -> some more pumping should be done during the next maintenance or whenever is possible;
  • the mSAS has been recentered in vacuum;
  • angular and position loops have been closed in vacuum;
  • the North arm has been re-aligned and re-locked in vacuum wihtout troubles.
Search Help